OpenStudy (anonymous):

The design of a digital box camera maximizes the volume while keeping the sum of the dimensions at 6 inches. If the length must be 1.5 times the height, what should each dimension be? Hint: Let x represent one of the dimensions, then define the other dimensions in terms of x. Help?

OpenStudy (anonymous):

I am assuming that it is a 2D shape. Is that right?

OpenStudy (anonymous):

Let the perimeter equal P P = 2l+2w 6 = 2*(1.5x)+2x 6 = 3x+2x 6 = 5x x = 6/5 x = 1.2 inches Width = 1.2 inches Height = 1.5*width = 1.8 inches

OpenStudy (anonymous):

thx!

OpenStudy (anonymous):

could you help me with another?

Directrix (directrix):

From Text-a-Friend @ridley9510 L = 1.5 H = (3/2) H (3/2)H + H + W = 6 so W = 6 - H - (3/2)H = 6 - (5/2) H V = L*H*W = (3/2) H * H * (6 - (5/2) H) volume is now a function of H alone and we can maximize it by taking the derivative. To make it easier, replace H by x. V = (3/2) x * x * (6 - (5/2) x) = 9 x^2 - (15/4) x^3 dV/dx = 18 x - (45/4) x^2 = 0 this has the solution x = 0 obtained by factoring out x. Then you have 18 - (45/4)x = 0 x = 72/45 as the other root. You can calculate the inflection points or simply graph out the function to be sure whether the points are minima or maxima. In this case, H = 72/45 = 1.6 is a maximum giving for the dimensions: H = 1.6 L = 2.4 W = 2.0 and the volume V = 7.68